OpenStudy (anonymous):

Gabe Amodeo, a nuclear physicist needs 60 liters of a 40% acid solution. He currently has a 20% solution and a 50% solution. How many liters of each class does he need to make the needed 60 liters of 40% acid solution. please help someone

OpenStudy (anonymous):

@spimentel Any idea?

OpenStudy (anonymous):

lol no,

OpenStudy (anonymous):

u?

OpenStudy (anonymous):

The goal is 60 L of 40%, so arrange it in the right side of the equation! Does it make some sense?

OpenStudy (anonymous):

no. :( can you show me step by step or solve the problem that way i can learn how to get the answer>

OpenStudy (anonymous):

I'm explaining step by step!

OpenStudy (anonymous):

I DONT KNOW HOW TO ARRANGE THE EQUATION

OpenStudy (chihiroasleaf):

I'm not sure that I fully understand the problem.., sorry... :( miss @ajprincess may can help ...:D

OpenStudy (anonymous):

OK THANK YOU.

OpenStudy (anonymous):

= 60 * 40% Done with the right side!

OpenStudy (anonymous):

Can you make a wild guess what to put on the LEFT side?

OpenStudy (anonymous):

60 LITERS *20%

OpenStudy (anonymous):

WRONG, because we already used 60L for the RIGHT side!

OpenStudy (anonymous):

What numbers left to arrange in the LEFT side?

OpenStudy (anonymous):

20 AND 50

OpenStudy (anonymous):

Yes, but we don't know the volume => 50x and 20 ( 60 - x)

OpenStudy (anonymous):

Can you write the complete equation now?

OpenStudy (anonymous):

(50-40):(40-20)=60

OpenStudy (anonymous):

Mixture means adding: 50x and 20 ( 60 - x) 50x + 20 ( 60 - x) = 40* 60 5x + 2 ( 60 - x) = 240 Now solve it!

OpenStudy (anonymous):

I GOT 40>

OpenStudy (anonymous):

It means 40L of 50% and...

OpenStudy (anonymous):

20L OF THE REMAINDING SOLUTION

OpenStudy (anonymous):

Yes, 20 L of 20%

OpenStudy (anonymous):

:) GOT IT. THANK SOOOOOOOOO MUCH, FOR BEING SOOOO PAT WITH ME

OpenStudy (anonymous):

Crystal clear :)

OpenStudy (anonymous):

YES LOL

OpenStudy (anonymous):

Practice makes perfect, then!

OpenStudy (anonymous):

DEFINITELY